# The Implication Problem for Functional Dependencies and Variants of Marginal Distribution Equivalences

### v1

@@ 3.3. Completeness and Armstrong relations
    Suppose that $\sigma=\mathop{=\!}(x)$. If $\sigma\in\Sigma$, then $x$ is determined by all the other variables in $\Sigma$. This means that $x$ is in the last component (the one with the highest scc-number), and thus the only value appearing in column $x$ is $k$, where $k$ is the number assigned to the maximal black clique of $x$. Hence, we have $\mathbb{X}\models\sigma$. If $\sigma\not\in\Sigma$, then either $x$ is not in the last component or there are no variables $y$ for which $\mathop{=\!}(y)\in\Sigma$. In either case, we have at some point added a tuple in which the position of $x$ is empty, and therefore there are more then one value appearing in column $x$, i.e., $\mathbb{X}\not\models\sigma$.
@@
    Since $\mathbb{X}$ is obtained by taking the uniform distribution over $X$, $\mathbb{X}\models x\approx^{*}y$ holds if and only if $|X(x)|=|X(y)|$. The latter happens if and only if $x$ and $y$ are in the same strongly connected component. Since any strongly connected component is itself the maximal blue clique, this is equivalent to $x$ and $y$ being in the same maximal blue clique in a strongly connected component, i.e., $x\approx^{*}y\in\Sigma$.
@@
    Since $\mathbb{X}$ is obtained by taking the uniform distribution over $X$, $\mathbb{X}\models x\approx y$ holds if and only if $X(x)=X(y)$. The latter happens if and only if $x$ and $y$ are in the same maximal black clique in a strongly connected component. This happens if and only if $x\approx y\in\Sigma$.
Example 3.5 ().
Let $\Delta=\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx x_{2},\mathop{=\!}(x_{2},x_{3}),x_{3}\approx^{*}x_{0},\mathop{=\!}(x_{4}),\mathop{=\!}(x_{5})\}$, and $\Sigma=\mathsf{cl}(\Delta)$. Then the strongly connected components of $G(\Sigma)$ are $\{x_{0},x_{1},x_{2},x_{3}\}$ and $\{x_{4},x_{5}\}$, the maximal red cliques are $\{x_{0},x_{1}\}$, $\{x_{2},x_{3}\}$, and $\{x_{4},x_{5}\}$, the maximal blue cliques are $\{x_{0},x_{1},x_{2},x_{3}\}$ and $\{x_{4},x_{5}\}$ and the maximal black cliques are $\{x_{0}\}$, $\{x_{1},x_{2}\}$, $\{x_{3}\}$, $\{x_{4}\}$, and $\{x_{5}\}$. The probabilistic team $\mathbb{X}$ for $\Sigma$ (as in Lemma 3.4) and the graph $G(\Sigma)$ are depicted im Table 2 and Figure 1.
Table 2. The probabilistic team $\mathbb{X}$.
$x_{0}$ $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $\mathbb{X}$
00 00 00 00 00 $1$ 1/4
00 00 $1$ $1$ 00 $1$ 1/4
$1$ $1$ 00 00 00 $1$ 1/4
$2$ $3$ $3$ $4$ 00 $1$ 1/4
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$
The construction of Lemma 3.4 does not yet give us Armstrong relations. This is because there might be nonunary functional dependencies $\sigma$ such that $\mathbb{X}\models\sigma$ even though $\sigma\not\in\Sigma$. In the lemma below, we show how to construct a probabilistic team for which $\mathbb{X}\models\sigma$ implies $\sigma\in\Sigma$ also for nonunary functional dependencies $\sigma$.
Lemma 3.6 ().
@@ 3.4. Non-existence of a complete $k$-ary axiomatization
Fix $k\in\{1,3,5,\dots\}$ and consider the set
$\Sigma_{k}=\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx^{*}x_{2},\dots,\mathop{=\!}(x_{k-1},x_{k}),x_{k}\approx^{*}x_{0}\},$
@@
$\Sigma=\Sigma_{k}$$\cup\{x_{i}\approx x_{i}\mid 0\leq i\leq k\}$
$\cup\{x_{i}\approx^{*}x_{i}\mid 0\leq i\leq k\}$
$\cup\{x_{i}\approx^{*}x_{j}\mid x_{j}\approx^{*}x_{i}\in\Sigma_{k}\}$
$\cup\{\mathop{=\!}(\bar{y},\bar{z})\mid\forall x_{i}\in\mathsf{var}(\bar{z}),\text{ if }x_{i}\not\in\mathsf{var}(\bar{y}),\text{ then }i\in\{1,3,\dots,k\}\text{ and }x_{i-1}\in\mathsf{var}(\bar{y})\}.$
@@
We claim that $\Sigma=\mathsf{cl}_{k}(\Sigma_{k})$. By the $k$-cycle rule (which is a $k+1$-ary axiom) and the completeness, we have $\Sigma_{k}\models\mathop{=\!}(x_{1},x_{0})$. Since $\mathop{=\!}(x_{1},x_{0})\not\in\Sigma$, it follows that there is no complete $k$-ary axiomatization for the FD+UMI+UMDE implication.
It is easy to check that $\Sigma\subseteq\mathsf{cl}_{k}(\Sigma_{k})$. Hence, it suffices to prove the following lemma:
Lemma 3.7 ().
@@
We prove the claim by induction on the length of the deduction.
@@
For the case $n=m+1$, let $\sigma\in\mathsf{cl}_{k}(\Sigma_{k})$ be such that the length of its deduction from $\Sigma_{k}$ is $n$. Let $\delta_{1},\dots,\delta_{k}\in\mathsf{cl}_{k}(\Sigma_{k})$ be the atoms that appear in the last step of deduction. (Note that it might be that $\delta_{i}=\delta_{j}$ for some $i\neq j$.) By the induction hypothesis $\delta_{1},\dots,\delta_{k}\in\Sigma$. Denote $\Delta:=\{\delta_{1},\dots,\delta_{k}\}$. Suppose then for a contradiction that $\sigma\not\in\Sigma$. We will show that then $\Delta\not\models\sigma$. Let $\delta\in\Sigma_{k}\backslash\Delta$. Note that by symmetry of the cycle, it suffices to consider the cases $\delta=\mathop{=\!}(x_{0},x_{1})$ and $\delta=x_{k}\approx^{*}x_{0}$.
@@
From Lemma 3.7, it follows that no sound $k$-ary axiomatization can be complete. Since any finite axiomatization would be $k$-ary for some natural number $k$, we obtain the following theorem:
Theorem 3.8 ().
@@ 4.1. Axioms for FD+UMI and FD+UIND
Axioms UMI1–UMI3, FD1–FD3, and the following variant of the cycle rule
If =($\text{If }\mathop{=\!}(x_{0},x_{1})\text{ and }x_{1}\approx x_{2}\text{ and }\dots\text{ and }\mathop{=\!}(x_{k-1},x_{k})\text{ and }x_{k}\approx x_{0},$
@@
for all $k\in\{1,3,5,\dots\}$, form a sound and complete axiomatization for the implication problem for FD+UMI. Soundness of this variant of the cycle rule immediately follows from the fact that $x\approx y$ implies $x\approx^{*}y$ and from the proof of the soundness of cycle rule in Section 3.2. For the completeness of the axiomatization, we can construct the Armstrong relations as in Section 3.3.
@@
For the unary functional dependencies and unary inclusion dependencies, we have the cycle rule:
If =($\text{If }\mathop{=\!}(x_{0},x_{1})\text{ and }x_{2}\subseteq x_{1}\text{ and }\dots\text{ and }\mathop{=\!}(x_{k-1},x_{k})\text{ and }x_{0}\subseteq x_{k},$
@@
for all $k\in\{1,3,5,\dots\}$.
@@ 4.2. Simulation
Theorem 4.1 ().
[…]
@@
  • $n=k+1$. The case of the reflexivity axiom “$x\subseteq x$” is clear.
@@
We begin by showing that the “from left to right” side of the claim holds.
@@
  • $n=k+1$. The case of the reflexivity axiom “$x\approx x$” is clear since “$x\subseteq x$” is an axiom in the other axiomatization.
@@
Next, we prove the “from right to left” side of the claim.
@@
  • $n=k+1$. The case of the reflexivity axiom “$x\subseteq x$” is clear since “$x\approx x$” is an axiom in the other axiomatization.
@@
    Suppose that the last axiom used in the deduction of $x\subseteq y$ is the cycle rule. Then $\Sigma^{*}\vdash\mathop{=\!}(x_{0},x_{1})\wedge x_{2}\subseteq x_{1}\wedge\dots\wedge\mathop{=\!}(x_{k-1},x_{k})\wedge x_{0}\subseteq x_{k}$, where $x\subseteq y\in\{x_{1}\subseteq x_{2},\dots,x_{k}\subseteq x_{0}\}$. By Lemma 4.2, we have $\Sigma^{*}\vdash x_{1}\subseteq x_{2}\wedge\dots\wedge x_{k}\subseteq x_{0}$, where the length of the deduction for each inclusion dependence is at most that of the corresponding reversed inclusion. This means that there is already a shorter deduction for $x\subseteq y$, and thus by the induction hypothesis $\Sigma\vdash x\approx y$.
@@
The above result clearly follows also from Theorem 5.1, but as seen, it is already a simple corollary of the simulation result.

### v2

@@ 3.2. Completeness and Armstrong relations
    Suppose that $\sigma=\mathop{=\!}(x)$. If $\sigma\in\Sigma$, then $x$ is determined by all the other variables in $\Sigma$. This means that $x$ is in the last component (the one with the highest scc-number), and thus the only value appearing in column $x$ is $k$, where $k$ is the number assigned to the maximal black clique of $x$. Hence, we have $\mathbb{X}\models\sigma$. If $\sigma\not\in\Sigma$, then either $x$ is not in the last component or there are no variables $y$ for which $\mathop{=\!}(y)\in\Sigma$. In either case, we have at some point added a tuple in which the position of $x$ is empty, and therefore there are more than one value appearing in column $x$, i.e., $\mathbb{X}\not\models\sigma$.
@@
    Suppose that $\sigma=x\approx^{*}y$. Since $\mathbb{X}$ is obtained by taking the uniform distribution over $X$, $\mathbb{X}\models x\approx^{*}y$ holds if and only if $|X(x)|=|X(y)|$. The latter happens if and only if $x$ and $y$ are in the same strongly connected component. Since any strongly connected component is itself the maximal blue clique, this is equivalent to $x$ and $y$ being in the same maximal blue clique in a strongly connected component, i.e., $x\approx^{*}y\in\Sigma$.
@@
    Suppose that $\sigma=x\approx y$. Since $\mathbb{X}$ is obtained by taking the uniform distribution over $X$, $\mathbb{X}\models x\approx y$ holds if and only if $X(x)=X(y)$. The latter happens if and only if $x$ and $y$ are in the same maximal black clique in a strongly connected component. This happens if and only if $x\approx y\in\Sigma$.
The construction of Definition 3.8 does not yet give us Armstrong relations. This is because there might be nonunary functional dependencies $\sigma$ such that $\mathbb{X}\models\sigma$ even though $\sigma\not\in\Sigma$, as demonstrated in the following example.
Example 3.11.
Let $\Delta=\{x_{0}\approx^{*}x_{1},x_{1}\approx^{*}x_{2}\}$ and $\Sigma=\mathsf{cl}(\Delta)$. The graph $G(\Sigma)$ and probabilistic team $\mathbb{X}$ for $G(\Sigma)$ constructed as in Definition 3.8 are depicted in Figure 2 and Table 4, respectively. Note that, e.g., $\mathbb{X}\models\mathop{=\!}(x_{0}x_{1},x_{2})$, but $\mathop{=\!}(x_{0}x_{1},x_{2})\not\in\Sigma$.
$x_{0}$$x_{1}$$x_{2}$
Figure 2. The graph $G(\Sigma)$ of Example 3.11.
Table 4. The probabilistic team $\mathbb{X}$ of Example 3.11 constructed as in Definition 3.8.
$x_{0}$ $x_{1}$ $x_{2}$ $\mathbb{X}$
00 00 00 1/4
00 $1$ $1$ 1/4
$1$ 00 $4$ 1/4
$2$ $3$ 00 1/4
In the lemma below, we show how to construct a probabilistic team for which $\mathbb{X}\models\sigma$ implies $\sigma\in\Sigma$ also for nonunary functional dependencies $\sigma$.
Lemma 3.12.
@@ 00 $1$ $1$ 1/4
Note that the size of the constructed Armstrong relations may be large with respect to the number of variables $|D|:=n$ because in the construction of Lemma 3.12, we may need to add up to $2^{n}-(n+2)$ new tuples to ensure that the nonunary FDs that are not implied by $\Sigma$ are not satisfied.
Example 3.13.
Let $\Sigma$ and $G(\Sigma)$ be as in Example 3.11. Table 5 depicts the probabilistic team $\mathbb{X}$ constructed as in Lemma 3.12. Note that now for all functional dependencies $\sigma$, $\mathbb{X}\models\sigma$ if and only if $\sigma\in\Sigma$.
Table 5. The probabilistic team $\mathbb{X}$ of Example 3.11 constructed as in Lemma 3.12.
$x_{0}$ $x_{1}$ $x_{2}$ $\mathbb{X}$
00 00 $1$ 1/7
00 $1$ 00 1/7
$1$ 00 00 1/7
00 00 00 1/7
00 $2$ $2$ 1/7
$2$ 00 $5$ 1/7
$3$ $4$ 00 1/7
@@ 3.3. Non-existence of a complete $k$-ary axiomatization
Fix $k\in\{3,5,\dots\}$ and consider the set
$\Sigma_{k}=\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx^{*}x_{2},\mathop{=\!}(x_{2},x_{3}),\dots,x_{k-2}\approx^{*}x_{k-1},\mathop{=\!}(x_{k-1},x_{k}),x_{k}\approx^{*}x_{0}\},$
@@
$\Sigma=\Sigma_{k}$$\cup\{x_{0}\approx x_{0},x_{1}\approx x_{1},\dots,x_{k}\approx x_{k}\}$
$\cup\{x_{0}\approx^{*}x_{0},x_{1}\approx^{*}x_{1},\dots,x_{k}\approx^{*}x_{k}\}$
$\cup\{x_{0}\approx^{*}x_{k},x_{k-1}\approx^{*}x_{k-2},\dots,x_{4}\approx^{*}x_{3},x_{2}\approx^{*}x_{1}\}$
$\cup\Sigma^{\prime},$
where $\Sigma^{\prime}$ is defined inductively as follows:
  • (i)
    If $\mathsf{var}(\bar{z})\subseteq\mathsf{var}(\bar{y})\subseteq\mathsf{var}(\bar{x})$, then $\mathop{=\!}(\bar{y},\bar{z})\in\Sigma^{\prime}$.
  • (ii)
    If $\mathop{=\!}(\bar{y},\bar{z})\in\Sigma^{\prime}$ and $i\in\{0,2,\dots,k-1\}$, then $\mathop{=\!}(\bar{y}x_{i},\bar{z}x_{i+1})\in\Sigma^{\prime}$.
  • (iii)
    If $\mathop{=\!}(\bar{y},\bar{z})\in\Sigma^{\prime}$, $\mathsf{var}(\bar{y})=\mathsf{var}(\bar{y}^{\prime})$, and $\mathsf{var}(\bar{z})=\mathsf{var}(\bar{z}^{\prime})$, then $\mathop{=\!}(\bar{y}^{\prime},\bar{z}^{\prime})\in\Sigma^{\prime}$.
@@
We claim that $\Sigma=\mathsf{cl}_{k}(\Sigma_{k})$. By the $k$-cycle rule (which is a $k+1$-ary axiom) and the completeness, we have $\Sigma_{k}\models\mathop{=\!}(x_{1},x_{0})$. By the definition of $\Sigma^{\prime}$, we have $\mathop{=\!}(x_{1},x_{0})\not\in\Sigma^{\prime}$. Since $\mathop{=\!}(x_{1},x_{0})\not\in\Sigma_{k}$, we have $\mathop{=\!}(x_{1},x_{0})\not\in\Sigma$. Hence, from $\Sigma=\mathsf{cl}_{k}(\Sigma_{k})$, it follows that there is no complete $k$-ary axiomatization for the FD+UMI+UMDE implication.
For $\Sigma=\mathsf{cl}_{k}(\Sigma_{k})$, we prove the following two lemmas:
Lemma 3.14.
$\Sigma\subseteq\mathsf{cl}_{k}(\Sigma_{k})$.
Proof.
Clearly $\Sigma_{k}\subseteq\mathsf{cl}_{k}(\Sigma_{k})$. The second, the third, and the fourth set of the union are subsets of $\mathsf{cl}_{k}(\Sigma_{k})$ because they are obtained from $\Sigma_{k}$ by using the rules UMI1, UMDE1, and UMDE2, respectively. Suppose then that $\sigma\in\Sigma^{\prime}$. The cases (i) and (iii) of the inductive definition of $\Sigma^{\prime}$ are clear by FD1 and FD2. Thus, we may assume that $\sigma=\mathop{=\!}(\bar{y}x_{i},\bar{z}x_{i+1})$ for some $i\in\{0,2,\dots,k-1\}$ and $\bar{y},\bar{z}$ such that $\mathop{=\!}(\bar{y},\bar{z})\in\mathsf{cl}_{k}(\Sigma_{k})$. As $\mathop{=\!}(x_{i},x_{i+1})\in\Sigma_{k}\subseteq\mathsf{cl}_{k}(\Sigma_{k})$, it suffices to show that for all $\bar{v},\bar{v}^{\prime},\bar{w},\bar{w}^{\prime}$, we can infer from the atoms $\mathop{=\!}(\bar{v},\bar{w})$ and $\mathop{=\!}(\bar{v}^{\prime},\bar{w}^{\prime})$, the atom $\mathop{=\!}(\bar{v}\bar{v}^{\prime},\bar{w}\bar{w}^{\prime})$ by using at most $k$-ary axioms. This can be done as follows: by applying FD3 to $\mathop{=\!}(\bar{v},\bar{w})$ and $\mathop{=\!}(\bar{v}^{\prime},\bar{w}^{\prime})$, we obtain $\mathop{=\!}(\bar{v}\bar{v}^{\prime},\bar{w}\bar{v}^{\prime})$ and $\mathop{=\!}(\bar{v}^{\prime}\bar{w},\bar{w}^{\prime}\bar{w})$. Then by applying FD1 to $\mathop{=\!}(\bar{v}^{\prime}\bar{w},\bar{w}^{\prime}\bar{w})$, we obtain $\mathop{=\!}(\bar{w}\bar{v}^{\prime},\bar{w}\bar{w}^{\prime})$. Then by applying FD2 to $\mathop{=\!}(\bar{v}\bar{v}^{\prime},\bar{w}\bar{v}^{\prime})$ and $\mathop{=\!}(\bar{w}\bar{v}^{\prime},\bar{w}\bar{w}^{\prime})$, we obtain $\mathop{=\!}(\bar{v}\bar{v}^{\prime},\bar{w}\bar{w}^{\prime})$, as wanted. ∎
Lemma 3.15.
@@
We prove the claim by induction on the length of the deduction of the atoms $\sigma\in\mathsf{cl}_{k}(\Sigma_{k})$.
@@
For the case $n=m+1$, let $\sigma\in\mathsf{cl}_{k}(\Sigma_{k})$ be such that the length of its deduction from $\Sigma_{k}$ is $n$. Let $\delta_{1},\dots,\delta_{k}\in\mathsf{cl}_{k}(\Sigma_{k})$ be the atoms that appear in the last step of deduction. (Note that it might be that $\delta_{i}=\delta_{j}$ for some $i\neq j$.) By the induction hypothesis $\delta_{1},\dots,\delta_{k}\in\Sigma$. Denote $\Delta:=\{\delta_{1},\dots,\delta_{k}\}$. Suppose then for a contradiction that $\sigma\not\in\Sigma$. We will show that then $\Delta\not\models\sigma$. Let $\delta\in\Sigma_{k}\backslash\Delta$. Such $\delta$ exists because $\Sigma_{k}$ has $k+1$ different atoms, whereas $\Delta$ only has at most $k$ atoms. Note that by symmetry of the cycle, it suffices to consider the cases $\delta=\mathop{=\!}(x_{0},x_{1})$ and $\delta=x_{k}\approx^{*}x_{0}$.
@@
From Lemmas 3.14 and 3.15, it follows that $\Sigma=\mathsf{cl}_{k}(\Sigma_{k})$, and therefore no sound $k$-ary axiomatization can be complete. Since any finite axiomatization would be $k$-ary for some natural number $k$, we obtain the following theorem:
Theorem 3.16.
@@ 4.1. Axioms for FD+UMI and FD+UIND
Axioms UMI1–UMI3, FD1–FD3, and the following variant of the $k$-cycle rules for all $k\in\{1,3,5,\dots\}$:
If =($\text{If }\mathop{=\!}(x_{i},x_{S_{k}(i)})\text{ for all even }0\leq i\leq k-1\text{ and }x_{j}\approx x_{S_{k}(j)}\text{ for all odd }1\leq j\leq k,$
@@
form a sound and complete axiomatization for the implication problem for FD+UMI. Soundness of this variant of the cycle rule immediately follows from the fact that $x\approx y$ implies $x\approx^{*}y$. For the completeness of the axiomatization, we can construct the Armstrong relations as in Section 3.2.
@@
For the unary functional dependencies and unary inclusion dependencies, we have the k-cycle rules for all $k\in\{1,3,5,\dots\}$:
If =($\text{If }\mathop{=\!}(x_{i},x_{S_{k}(i)})\text{ for all even }0\leq i\leq k-1\text{ and }x_{S_{k}(j)}\subseteq x_{j}\text{ for all odd }1\leq j\leq k$
@@
The soundness of $k$-cycle rules for unary functional dependencies and unary inclusion dependencies was shown in (Cosmadakis et al., 1990) but we include a brief proof for the sake of self-containment.
Suppose now that the antecedent of the $k$-cycle rule holds for $\mathbb{X}$. It then follows that
$|X(x_{0})|\geq|X(x_{1})|\geq|X(x_{2})|\geq\dots\geq|X(x_{k-1})|\geq|X(x_{k})|\geq|X(x_{0})|,$
which implies that $|X(x_{i})|=|X(x_{j})|$ for all $i,j\in\{0,\dots k\}$. If $i,j$ are such that $\mathbb{X}\models\mathop{=\!}(x_{i},x_{j})$, then $\mathbb{X}\models\mathop{=\!}(x_{j},x_{i})$ follows from the same argument as in the case of $k$-cycle rules for unary functional dependencies and unary marginal distribution equivalence in Section 3.1. If $i,j$ are such that $\mathbb{X}\models x_{i}\subseteq x_{j}$, then $X(x_{i})\subseteq X(x_{j})$. When combined with the fact that $|X(x_{i})|=|X(x_{j})|$, we obtain that $X(x_{j})\subseteq X(x_{i})$, i.e., $\mathbb{X}\models x_{j}\subseteq x_{i}$.
@@ 4.2. Simulation
Theorem 4.1.
Let $\Sigma_{\text{FD}}$ and $\Sigma_{\text{UMI}}$ be sets of functional dependencies and unary marginal identities, respectively. Denote $\Sigma=\Sigma_{\text{FD}}\cup\Sigma_{\text{UMI}}$, and define $\Sigma^{*}=\Sigma_{\text{FD}}\cup\{x\subseteq y\mid x\approx y\in\Sigma_{\text{UMI}}\text{ or }y\approx x\in\Sigma_{\text{UMI}}\}$. Then for all functional dependencies $\sigma$ and unary marginal identities $x\approx y$
@@
  • $n=m+1$. The case of the reflexivity axiom “$x\subseteq x$” is clear.
@@
We begin by showing that the “from left to right” sides of the two items in the theorem hold.
@@
  • $n=m+1$. The case of the reflexivity axiom “$x\approx x$” is clear since “$x\subseteq x$” is an axiom in the other axiomatization.
@@
Next, we prove the “from right to left” sides of the items.
@@
  • $n=m+1$. The case of the reflexivity axiom “$x\subseteq x$” is clear since “$x\approx x$” is an axiom in the other axiomatization.
@@
    Suppose that the last axiom used in the deduction of $x\subseteq y$ is the cycle rule. Then $\Sigma^{*}\vdash\mathop{=\!}(x_{0},x_{1})\wedge x_{2}\subseteq x_{1}\wedge\dots\wedge\mathop{=\!}(x_{k-1},x_{k})\wedge x_{0}\subseteq x_{k}$, where $x\subseteq y\in\{x_{1}\subseteq x_{2},\dots,x_{k}\subseteq x_{0}\}$. By Lemma 4.2, we have $\Sigma^{*}\vdash x_{1}\subseteq x_{2}\wedge\dots\wedge x_{k}\subseteq x_{0}$, where the length of the deduction for each inclusion dependence is at most that of the corresponding reversed inclusion, i.e., at most of the length $m$. Since $y\subseteq x$ already appeared in the antecedent of the used rule, this means that there is already a shorter deduction for $x\subseteq y$ (the one obtained from Lemma 4.2), and thus by the induction hypothesis $\Sigma\vdash x\approx y$.
@@
The above result clearly follows also from Theorem 5.1 to be proved in the next section, but as seen, it is already a simple corollary of the simulation result.